# Masked Language Models are Good Heterogeneous Graph Generalizers

### v1

@@ Appendix E Impact of Different LMs and LLMs
The impact of different LMs is worth investigating. Considering the ongoing development of the causal language modeling (CLM) Achiam et al. (2023); Touvron et al. (2023), we not only evaluate MLM-based models, such as RoBERTa Liu (2019) and DistilBERT Sanh (2019), but also explore the performance of LLaMA Touvron et al. (2023) (specifically, LLaMA-3.1-8B-Instruct) in our experiments. To ensure a fair comparison, we leverage instructions to restrict LLaMA’s predictions to the constrained target vocabulary. We also report the training time per batch for each model, with a batch size of 16. For MLM-based language models, full parameter fine-tuning are applied, while LLaMA is fine-tuned using LoRA Hu et al. (2021).
@@ Appendix F Hyperparameter Study
We evaluated the impact of two hyperparameters: maximum token sequence length (Max Length) and fine-tuning epochs on MLM4HG performance. As shown in Figure 3, increasing the token sequence length enhances both node classification and link prediction, indicating that longer metapath-based sequences help LM capture more structural and semantic information. Additionally, as fine-tuning epochs increase, both node classification and link prediction tasks exhibit initial improvement followed by stabilization. Thus, we set the fine-tuning epochs to 3 in our experiments.
@@ Appendix G Zero-shot Link Prediction Evaluation Using AUC and AP
Appendix H Case Study
@@
Appendix I Examples and Parsing Results

### v2

@@ Appendix E Impact of Different LMs and LLMs
The impact of different LMs is worth investigating. Considering the recent advances of causal language modeling (CLM) approachesAchiam et al. (2023); Touvron et al. (2023), we not only evaluate MLM-based models, such as RoBERTa Liu (2019) and DistilRoBERTa Sanh (2019), but also explore the performance of LLaMA3 Touvron et al. (2023) (specifically, LLaMA-3.1-8B-Instruct) and Qwen3 Yang et al. (2025) (specifically, Qwen3-8B) in our experiments. To ensure a fair comparison, we leverage instructions to restrict CLM-based models’ predictions to the constrained target vocabulary. We also report the training time per batch for each model, with a batch size of 16. For MLM-based language models, full parameter fine-tuning are applied, while CLM-based models are fine-tuned using LoRA Hu et al. (2021).
@@ Appendix F Hyperparameter Study
We conducted a detailed evaluation of two key hyperparameters: maximum token sequence length (Max Length) and the number of fine-tuning epochs (Fine-tuning epochs), to understand their influence on the performance of MLM4HG. As illustrated in Figure 3, extending the token sequence length consistently improves results for both node classification and link prediction tasks. This suggests that longer metapath-based sequences enable the language model to capture richer structural and semantic context from the heterogeneous graph. Furthermore, increasing the number of fine-tuning epochs initially leads to noticeable performance gains for both tasks, after which the results tend to stabilize. Based on this observation, we fixed the number of fine-tuning epochs to 3 in all reported experiments to balance performance and training efficiency.
@@ Appendix G Zero-shot Link Prediction Evaluation Using AUC and AP
Appendix H Limitations and Broader Impacts
We discuss the limitations and broader impacts of MLM4HG in this section. Although our exploratory experiments on metapath-based sequences in Section 4.3 demonstrate that MLM4HG is able to extract and leverage structural signals from such sequences for downstream tasks, the semantic information encoded in textual content remains crucial for language models. Therefore, applying LMs to non-textual graph data still requires further exploration. Moreover, additional investigation is needed to assess the model’s generalizability across a wider range of graph tasks.
MLM4HG is a simple yet effective framework for generalizable heterogeneous graph learning. Its potential impact extends to a variety of real-world applications involving HGs, such as social network analysis, knowledge representation, and recommender systems. While the ability to generalize across domains and tasks can enhance the development of robust AI systems, it may also introduce concerns. For example, the model could be misused in illegal activities, leading to unintended consequences such as the perpetuation of biases or the misuse of predictive insights. We encourage future work to focus on enhancing the transparency and fairness of graph language models, especially in sensitive applications.
Appendix I Case Study
@@
Appendix J Examples and Parsing Results